OpenStudy (anonymous):

find the trig value below without a calculator. sec theta=7/4, find cos theta+cos theta tan^2 theta

OpenStudy (anonymous):

\[\sec \theta=\frac{ 7 }{ 4 }, \cos \theta +\cos \theta \tan ^{2} \theta\]

OpenStudy (anonymous):

okay I know that sec theta=sec(- theta)

OpenStudy (anonymous):

I also know that sec= 1/cos

OpenStudy (anonymous):

but i cant put it together.

OpenStudy (anonymous):

any luck @NoelGreco ?

OpenStudy (noelgreco):



Directrix (directrix):

@natnatwebb cos theta+cos theta tan^2 theta Factor out cos(theta) to get cos(theta)[ 1 + tan^2 (theta)]=

Directrix (directrix):

[1 + tan^2 (theta)]= is a trig identity. Look it up if you don't remember it at the moment.

Directrix (directrix):

You'll also need the identity cos(theta) = 1/sec(theta) and this answer will fall right out.

OpenStudy (anonymous):

okay, every single question i did by myself this afternoon i got wrong.

OpenStudy (anonymous):

okay...lemme think.

Directrix (directrix):

Just take one question at the time.

OpenStudy (anonymous):

okay, how did you factor out the cos?

OpenStudy (anonymous):

and and...the 1, where did that sucker come from

OpenStudy (anonymous):